For 80% of lectures, Professor John Smith arrives on time and starts lecturing with delay T. When Professor is late, the starting time delay T is uniformly distributed between 0 and 300 seconds. Find the probability distiribution and probability density function of random variable T.



1
Expert's answer
2021-01-04T18:58:28-0500

For continuous uniform distribution probability density function will be:

"f(T) = \begin{cases}
 1/(b-a) &\text{for } a\le T\le b \\
 0 &\text{for T<a or T>b} 
\end{cases}"

"f(T) = \begin{cases}
 1/300 &\text{for } 0\le T\le 300 \\
 0 &\text{for } T>300
\end{cases}"

Let's describe probability distribution:

P(0"\le"T"\le"100) = "\int_0^{100}1/300 dx=100/300=1/3"

P(100"\le"T"\le"200) ="\int_{100}^{200}1/300 dx=(200-100)/300=1/3"

P(200"\le"T"\le"300) = "\int_{200}^{300}1/300 dx=(300-200)/300=1/3"

P(T>300) = 0